3 rmed in 3a4 s e Co s easy to show th EelxJ-EY

Do all parts.
Answer #1

(a) FY(t) = P(y < t) = P(c-X < t) = P(X > c-t) = 1-Fx(c-t) Fx(x) = 0 C Now, take Y = X2 then fy(t) = 0 otherwise Since fx(x)-